Melanie can paint a certain barn by herself in xday. her helper Malissa can paint the same barn by herself in 2x days. Write a rational expression for the fraction of the barn that they complete in one day by working together. Evaluate the expression for x=5.
Answer by stanbon(75887)

Melanie can paint a certain barn by herself in xday. her helper Malissa can paint the same barn by herself in 2x days. Write a rational expression for the fraction of the barn that they complete in one day by working together. Evaluate the expression for x=5.
------------
Melanie DATA;
Time = x days/job ; Rate = 1/x job/day
---------------------
Melissa DATA:
Time = 2x days/job ; Rate = 1/2x job/day
---------------------
Together data:
Time = k days/job ; Rate = 1/k job/day
--------------------------------------------
EQUATION:
rate + rate = togther rate
1/x + 1/2x = 1/k
3x/2x^2 = 1/k
k = 2x^2/3x = (2/3)x
============================
If x = 5 the two will require (2/3)5 = 10/3 = 3 1/3 days to do the job.
==========================
Cheers,
Stan H.